# Multiple-Resolution Tokenization for Time Series Forecasting with an Application to Pricing

### v1

@@ 4 Application to a Forecasting Problem in Pricing
The data for the experiments is real markdown data from a very large European retailer. The markdown pricing prediction problem is to estimate the sales of an item given a series of prices and other contextual information. This prediction problem is understood as particularly difficult. The theory is that each time there is a price reduction we observe a spike in sales which then decays. Subsequent price reductions produce a similar spike and decay effect which typically reduces in magnitude with each reduction. This highly non-linear behaviour is difficult to model with simple approaches. This is further exacerbated by the high levels of noise associated with consumer behaviour owing to stochastic factors like the weather or competitor behaviour and the human factor in actually delivering the pricing strategy in stores. Historical data is limited to a narrow set of tried and tested pricing strategies such as $25\%$ off then $50\%$ then $75\%$ adding to the epistemic uncertainty. The downstream decision problem of setting reduced prices for items that are about to expire sees about a million decisions made every day.
@@
In the case of markdowns the inclusion of auxiliary data into the model is necessary as it is otherwise not possible to use it as a hypothesis generator for new pricing strategies. The decision variable (price) must be included as an input in the prediction model in order to be able to compare predictions under different markdown strategies. We try to predict both the full-price and the reduced price-sales concurrently, this defines the channels in $C$. This reflects the operational characteristics of the retailer where some proportion items that are about to expire get moved to a discounted section.
One view of the utility of auxiliary information beyond the necessary decision variables is that it enables the aggregation of a larger number of time series in a single training corpus with more learning potential. We can learn tokenizations which capture the relatively specific context of each time series, which leads to a more general purpose prediction model for sales. In the case of markdowns individual time series are relatively short and highly contextual. We take each series to be defined by as the unique combination of product, store, and expiration date. We want to aggregate all products of interest across all shops for all dates in a single model. We include auxiliary information because we do not believe that there is sufficient information in the starting course of a series to effectively differentiate amongst products and stores and expiration dates. We use following auxiliary information:
@@
Normalisation
Data Setting
The main issue encountered when attempting to apply Multiple Resolution Tokenisation to pricing data is that the data has a distinct real-world flavour. Sales are aggregated within varying blocks of time that represent a duration of displayed price. The time series is irregularly spaced and each series has very few observations (at most 4). The data does not contain time series of sales when markdowns are not present. Each set of scan observations must be treated as its own series. Since only markdowns are observed, there is a clear selection effect as the process to select markdowns is biased toward existing more profitable strategies. This selection bias in the data generating process may result in poorer generalisation and overconfidence. The signal is also somewhat limited as there is only a narrow scope of markdown strategies that are regularly implemented. The dataset features many missing values and heuristics to deal with data acquisition problems.
4.1 Preprocessing
From a model architecture standpoint this dataset poses two key problems: irregularly spaced sampling, and the varying length of the series.
@@ 4.1 Preprocessing
4.2 Experiments
The experiments were defined by the range of year-weeks they contained. The experiments are limited to just one week due to memory constraints. Each experiment contains the markdown data for one week across all shops and all items. For any experiment the series corresponding to the last $20\%$ (rounded up) of the due-dates were held out as the testing set, the $15\%$ (rounded up) of due-dates immediately preceding were used as the validation set and the remaining as the training set.
@@ 4.2 Experiments
Results
Table 1: Experiment week 30 with $l=28$.
Model MSE MAE full price MSE full price MAE
internal NaN NaN 0.0821 0.1443
MRP 0.0717 0.1323 0.0895 0.1447
PatchTST 0.0765 0.1349 0.0963 0.1480
DLinear 0.0742 0.1380 0.0959 0.1549
Table 2: Experiment week 45 with $l=37$.
Model MSE MAE full price MSE full price MAE
internal NaN NaN 0.0764 0.1214
MRP 0.0269 0.0661 0.0474 0.0877
PatchTST 0.0660 0.1190 0.0926 0.1370
DLinear 0.0521 0.1156 0.0745 0.1344
The model is competitive with internal predictions and an improvement on existing models in the week $30$ experiment and significantly outperforms all other models on week $45$. Note that the test scores have been reverted to their original scales. Given the complexity of the prediction task across a range of products and locations and the necessary adjustments to fit a sequence-to-sequence model the result for week $45$ provides strong evidence that MRT is a powerful model which could deliver much with hyperparameter optimisation and scaling. The week ranges on these experiments are somewhat limited but we are confident that over longer ranges these models would be even more competitive due to stronger auxiliary embeddings (especially for global temporal information) and general scaling effects.
One contrarian explanation of week 45 results is that many values are 0 and our network learns to predict near 0s in many cases. The many values being 0 case is somewhat credible given our selection of only the longest series; the longer duration could reflect that getting rid of stock in those cases is difficult. An encouraging sign is that the predictions for reduced price sales are better than those for the full price sales (assuming similar scales across channels). Since we have shown that we are competitive or better than internal models for full price sales, the reduced price sales predictions show even more promise.
4.3 Discussion
The experiments are somewhat limited in scope. This is mainly due to computational constraints and the delayed timing of the project. More experimentation is needed to weigh the evidence on this architecture. With more compute the experiments could be extended to much larger dataset ranges and latent spaces. Scaling properties are clearly of interest. Ablation would demonstrate what relative benefit there is for each of our tokenisation procedures. Hyperparameter exploration in the space of resolution sets would help determine heuristics on what resolution combinations are performant and uncover good heuristics like perhaps including multiple copies of the same resolution if it is highly involved in attention patterns.
Due to commercial sensitivity the dataset cannot be shared [^2] We provide the model code here.. We acknowledge that this is a serious limitation from a reproducibility standpoint, but believe that the experiments performed here are a substantial contribution to research on transformers for forecasting. This is because most existing architectures have been empirically evaluated on a narrow set of datasets which are typically small scale such as ETT (Zhou et al., 2021). What is most likely dataset bias has caused a shift of interest away from using auxiliary or cross-series information. In contrast we demonstrate that our transformer architecture outperforms currently used in-house methods, DLinear, and PatchTST on a messy real-world problem. The work presents evidence that for certain forecasting problems modelling auxiliary data and cross series information adds to generalisation. More ablation is needed to understand what aspects of our tokenisation strategies are the most important in supporting this claim.
5 Conclusion
The experiments provide clear evidence that our proposed MRT architecture works well on a noisy real-world problem. This is achieved with an architecture which includes both auxiliary variables and models cross-series information which otherwise is often avoided in transformer architectures for time series forecasting. The modules designed for this purpose are novel as is the practice of treating all of their outputs as tokens in a single attention mechanism. While our experimentation is limited one of the experiments shows a substantial improvement in performance over existing architectures. Given the difficulty of the underlying prediction problem this is a significant observation in favour of testing and potentially deploying transformer architectures on difficult real-world forecasting problems in the pricing domain and beyond.

### v2

@@ 5 Application to a Forecasting Problem in Pricing
5.1 Real Markdown Problem
Markdowns are a subset of pricing problems (Phillips, 2021) concerned with reducing prices to stock that can no longer be sold in a store. Reduced to clear process is commonly applied to perishable items that are due to expire, to products before the end of their seasonal use or to vacate space for new product lines. The decision problem is to set a series of price reductions such that an objective is optimised (Aleksandar Kolev & Arafailova, 2023). The objective is typically revenue combined with some sustainability based measure. A key aspect of such problems is predicting the sales of an item in response to a set price in a given context. A forecasting model which incorporates price as an auxliary variable can be effectively be utilised as a simulator to narrow down pricing strategies for empirical testing. Automated Markdown strategies, using advanced forecasting and optimisation techniques, have been develop by many large merchandisers including Zara (Caro & Gallien, 2012), ASOS (Loh et al., 2022), Walmart (Chen et al., 2021), and Tesco (Kolev et al., 2023). These strategies have been applied to both online and physical stores and have shown a general increase of revenue with a reduction of unsold items by the end of the markdown window. There are considerable financial, environmental, operational and reputational risk in applying a new pricing strategy at scale. A good simulator enables a virtual exploration of the pricing strategy hypothesis space which can help select potentially more profitable and sustainable pricing strategies for testing.
@@ 5.1 Real Markdown Problem
The data for the experiments is real markdown data from one of the largest European grocery retailers. The markdown pricing prediction problem is to estimate the sales of an item given a series of past sales, prices and other contextual information. This prediction problem is understood as particularly difficult. The theory is that each time there is a price reduction we observe a spike in sales which then decays (Phillips, 2021). Subsequent price reductions produce increasingly dampened spiking. This highly non-linear behaviour is difficult to model with simple approaches. Walmart’s and probably many others (Chen et al., 2021) first algorithmic generated reduced to clear prices have performed worse than standard strategies. The markdown optimisation problem is further exacerbated by the high levels of noise associated with consumer behaviour owing to stochastic factors like the weather or competitor behaviour. The product allocation and price application is usually a manual process which adds the human factor to the data generation process. Historical data is limited to a narrow set of tried and tested pricing strategies such as $25\%$ off then $50\%$ then $75\%$ adding to the epistemic uncertainty. A typical large retailer’s downstream decision problem of setting reduced prices for items that are about to expire faces millions of decisions made every day.
We try to predict both the full-price and the reduced price-sales concurrently, this defines the channels in $C$. Both of these predictions are crucial for finding the optimal revenue and waste management as any errors of the full-price models propagate further to an increased number of products that appear in the reduced to clear process. We include more details about the experiment in the appendix.
5.1.1 Experiments
The experiments were defined by the range of year-weeks they contained. The experiments are limited to just one week due to memory constraints. Each experiment contains the markdown data for one week across all shops and all items. For any experiment the series corresponding to the last $20\%$ (rounded up) of the due-dates were held out as the testing set, the $15\%$ (rounded up) of due-dates immediately preceding were used as the validation set and the remaining as the training set. We set our forecast horizon to $f=24$ and adjust the lookback horizon accordingly with the longest sequence in a given experiment. We note that the model produces $52$ tokens.
Table 1: Test results for experiment week 30 with $l=28$.
Model MSE MAE full price MSE full price MAE
internal NaN NaN 0.0821 0.1443
MRP 0.0717 0.1323 0.0895 0.1447
PatchTST 0.0765 0.1349 0.0963 0.1480
DLinear 0.0742 0.1380 0.0959 0.1549
Table 2: Test results for experiment week 45 with $l=37$.
Model MSE MAE full price MSE full price MAE
internal NaN NaN 0.0764 0.1214
MRP 0.0269 0.0661 0.0474 0.0877
PatchTST 0.0660 0.1190 0.0926 0.1370
DLinear 0.0521 0.1156 0.0745 0.1344
The model is competitive with internal predictions and an improvement on existing models in the week $30$ experiment and significantly outperforms all other models on week $45$. Note that the test scores have been reverted to their original scales. Given the complexity of the prediction task across a range of products and locations and the necessary adjustments to fit a sequence-to-sequence model the result for week $45$ provides strong evidence that MRT is a powerful model which could deliver much with hyperparameter optimisation and scaling. The week ranges on these experiments are somewhat limited but we are confident that over longer ranges these models would be even more competitive due to stronger auxiliary embeddings and general scaling effects.
One contrarian explanation of week 45 results is that many values are 0 and our network learns to predict near 0s in many cases. The many values being 0 case is somewhat credible that the dataset only contains longer series; the longer duration could reflect that getting rid of stock in those cases is difficult. An encouraging sign is that the predictions for reduced price sales are better than those for the full price sales (assuming similar scales across channels). Since we have shown that we are competitive or better than internal models for full price sales, the reduced price sales predictions show even more promise.
5.2 Favorita Dataset Problem
In addition to experiments with private data we replicated a markdown type forecasting problem on a public dataset. We use the Favorita dataset (Corporación Favorita, 2017) which contains the daily sales of a range of items across a range of locations from a South American grocer. In addition it includes auxiliary information such as whether an item was under promotion which is somewhat analogous to the markdown problem. We narrow the dataset to only include the 78 items present in all locations and sold on all dates. We set $f=16$ and $l=32$ and apply a sliding window approach to create unique series. We select locations as our channels. This means that for each input we simultaneously predict the sales for the next 16 days for a given item in all 53 shops. The implied assumption from a cross-series perspective is that we believe that there may be leading indicators in some locations for other locations. We use the same holdout approach and base hyperparameters as in the first experiment unless noted otherwise. We conducted three sets of experiments to ablate the model and empirically establish the usefulness of modules.
5.2.1 Resolution Set Sizes
Picking a resolution set involves selecting from exponentially many options. We restrict possible resolution sets to including any nonempty combination of $k\in\{1,2,4,8,16\}$, so 31 possible combinations. We train each model once due to the large number of experiments and acknowledge that there is a significant variance in outcomes. Note that both TVKT and CST were included in these experiments. The results are presented in figure 5.
Figure 5: Favorita experiment test results for the resolution set ablation.
We find that the best performing configurations are ones with more than one resolution, but that more resolutions does not necessarily mean improved performance. The optimal resolution set for a problem depends on the structure of the underlying process. Interestingly the best performing single resolution model is with $k=1$, which means extracting a single token to represent all past data. The results here strongly support using multiple resolutions, but that the choice of resolution requires careful selection. We pick the second best performing resolution set $\{1,2,4,16\}$ as a baseline in the following ablations.
5.2.2 Module Inclusion
We test four variants of the model: no auxiliary variable tokens, just TVKT, just CST, and both groups. In addition we test the two models we compared MRT with in the previous experiment. We run each experiment five times and display the results with standard deviations in table 3. The best performing performance on average was achieved by our MRT model without additional tokens. This further supports the use of MRT, though the gain in performance is slim in exchange for about an order of magnitude increase in computation when comparing it to PatchTST. As expected, including auxiliary variable tokens induces a lot of variance in results due to potential overfitting, but expect that this could be reduced by balancing the number of auxiliary tokens against MRP tokens. We note the best performance within each ablation and that overall the best performance in any of these experiments was achieved when including TVKT which is encouraging as they have to be included to be useful for pricing problems. CST appear to worsen performance which signals that there is little persistent leading information across channels for this problem. We exclude just using CST for the next ablation.
Table 3: Favorita experiment test results for module inclusion ablation. RMSE loss.
Measure None PatchTST TVKT CST Both DLinear
mean 16.783 16.823 17.281 17.376 17.621 19.192
std 0.038 0.047 0.751 0.185 1.294 0.320
min 16.756 16.771 16.614 17.145 16.977 18.804
5.2.3 Scaling
We ran a small scaling experiment with four levels of scaling $d_{m}\in\{32,64,96,128\}$ and three module combinations: none, TVKT, and both. We avoided larger scales due to the small size of the dataset (approximately 17k series). The rule for other parameters was $d_{ff}=2d_{m}$, the latent size in the mixer for CST is set as $d_{cross}=\frac{1}{4}d_{m}$, number of heads was selected as $\frac{1}{8}d_{m}$ and all models had two transformer blocks. The results are displayed in table 4. Interestingly TVKT performs better at a smaller latent size which may be explained by the size of the dataset. The small scale TVKT is also the best performing model of all in this experiment, tough this may not be consistently true due to previously observed variance in training models with auxiliary tokens. The other two ablations appear to scale with model size, though this effect for including both could be explained by gains in performance for MRP tokens as the ablation without any auxiliary information marginally improves with performance with scale.
Table 4: Favorita experiment test results for scaling $d_{m}$ ablation. RMSE loss.
$d_{m}$ Both TVKT None
32 17.462 16.555 16.847
64 17.431 16.803 16.720
96 16.848 17.034 16.680
128 16.884 17.056 16.646
5.3 Discussion
On the private dataset we demonstrate that our transformer architecture outperforms currently used in-house methods, DLinear, and PatchTST on a messy real-world problem. We ablate the model on a forecasting problem derived from a public dataset and demonstrate that our architecture is competitive with PatchTST and has better best-cases. We provide evidence that multiple resolution modelling outperforms single resolution patching. The experiments demonstrate that for certain forecasting problems modelling auxiliary data can add to generalisation contrary to common research trends, specifically the addition of TVKT which ended up producing the best performing model of all. More experimentation is needed to demonstrate the utility of the channel mixer and our suspicion is that it could be of value for the right channel structure.
6 Conclusion
The experiments provide clear evidence that our proposed MRT architecture works well on a noisy real-world problem. This is achieved with an architecture which includes both auxiliary variables which otherwise is often avoided in transformer architectures for time series forecasting. The modules designed for this purpose are novel as is the practice of passing all of their output tokens to a single attention mechanism. Our ablations show the fascinating effects different resolution sets have on performance and provide evidence in favour of multiple resolution approaches. One of the experiments shows a substantial improvement in performance over existing architectures. Given the difficulty of the underlying prediction problem this is a significant observation in favour of testing and potentially deploying transformer architectures on difficult real-world forecasting problems in the pricing domain and beyond.
Appendix A Additional Details About Experiments
A.1 Real Markdown Problem
Data Setting
The main issue encountered when attempting to apply Multiple Resolution Tokenisation to pricing data is that the data has a distinct real-world flavour. Sales are aggregated within varying blocks of time that represent a duration of displayed price. The time series is irregularly spaced and each series has very few observations (at most 4). The data does not contain time series of sales when markdowns are not present. Each set of scan observations must be treated as its own series. Since only markdowns are observed, there is a clear selection effect as the process to select markdowns is biased toward existing more profitable strategies. This selection bias in the data generating process may result in poorer generalisation and overconfidence. The signal is also somewhat limited as there is only a narrow scope of markdown strategies that are regularly implemented. The dataset features many missing values and heuristics as a result of data acquisition problems.
@@ A.1 Real Markdown Problem
In the case of markdowns the inclusion of auxiliary data into the model is necessary as it is otherwise not possible to use it as a hypothesis generator for new pricing strategies. The decision variable (price) must be included as an input in the prediction model in order to be able to compare predictions under different markdown strategies. We use only two channels: full price and reduced price sales which we attempt to forecast simultaneously. This reflects the operational characteristics of the retailer where some proportion items that are about to expire get moved to a discounted section.
In the case of markdowns individual time series are relatively short and highly contextual. We take each series to be defined by as the unique combination of product, store, and expiration date. We want to aggregate all products of interest across all shops for all dates in a single model. We include auxiliary information because we do not believe that there is sufficient information in the starting course of a series to effectively differentiate amongst products and stores and expiration dates. We use following auxiliary information:
@@
A.1.1 Preprocessing
From a model architecture standpoint the real markdown problem dataset poses two key problems: irregularly spaced sampling, and the varying length of the series.
@@ A.1.1 Preprocessing
Normalisation
A.1.2 Model details
Note that we use the same setting in the second experiment unless otherwise noted.
@@ A.1.2 Model details
A.2 Favorita
Auxiliary Information
We use following auxiliary information:
  • Static:
    • Global: item class, item family, item id, is perishable
    • Specific: state, city, store id
  • TVK:
    • Global: numerical distance to holidays, day of week, day of month, month
    • Specific: on promotion (analogous to pricing variable)